# Improved subsample-and-aggregate via the private modified winsorized mean

### v1

@@ Abstract
We develop a univariate, differentially private mean estimator, called the private modified winsorized mean, designed to be used as the aggregator in subsample-and-aggregate. We demonstrate, via real data analysis, that common differentially private multivariate mean estimators may not perform well as the aggregator, even with a dataset with 8000 observations, motivating our developments. We show that the modified winsorized mean is minimax optimal for several, large classes of distributions, even under adversarial contamination. We also demonstrate that, empirically, the modified winsorized mean performs well compared to other private mean estimates. We consider the modified winsorized mean as the aggregator in subsample-and-aggregate, deriving a finite sample deviations bound for a subsample-and-aggregate estimate generated with the new aggregator. This result yields two important insights: (i) the optimal choice of subsamples depends on the bias of the estimator computed on the subsamples, and (ii) the rate of convergence of the subsample-and-aggregate estimator depends on the robustness of the estimator computed on the subsamples.
@@ 1 Introduction
Subsample-and-aggregate is a powerful algorithm that can produce a differentially privately version of any statistic [29]. For instance, through subsample-and-aggregate, complex machine learning models can be made private easily, with minimal additional programming. The algorithm is a two-step procedure, relying on a “subsample” step and an “aggregate” step. The idea, in brief, is to first split the data into disjoint subsets. Then, on each of these subsets, a nonprivate version of the desired statistic is computed. Next, the set of nonprivate statistics are combined, using a differentially private aggregator, producing a private estimate of the parameter of interest. Generally, if the mean of the subsamples is desired, the private aggregator is a basic differentially private clipped mean [29] or a widened winsorized mean [33]. However, to achieve pure differential privacy, both of these aggregators require tight input bounds on the statistic.
It is natural to consider the more recently developed, differentially private mean estimators, which have a weaker dependence, if any, on input bounds, as the private aggregator. However, currently, replacing the aggregator with any of the existing differentially private mean estimators comes with at least one of the following drawbacks: (i) only a weaker notion of differential privacy is satisfied, (ii) many subsamples are required for a stable mean estimate, resulting in large sample size requirements, (iii) the new algorithm has strong dependence on tuning parameters, (iv) strong assumptions on, or knowledge of, the population distribution is required, or (v) the estimator is not practically computable.
Here, we develop a new, univariate, differentially private mean estimator, called the private modified winsorized mean. Our estimator is designed to be used as the aggregator in subsample-and-aggregate and avoids these drawbacks. The private modified winsorized mean can satisfy either pure or zero-concentrated differential privacy, performs well in small samples or under adversarial contamination, has weak dependence on input bounds and other tuning parameters, and can be used to estimate the mean of a general distribution, where only two moments need to exist. The proposed estimator combines the winsorized mean of Lugosi and Mendelson, [28] and the private quantile estimation procedure of Durfee, [13].
@@ 1.1 Related work
Differentially private mean estimation has been well-studied. In the univariate setting, clipped and trimmed means have been the main approach [18, 24, 9, 23, 34, 13]. The clipped mean procedure operates by selecting an interval $(a,b)$, either empirically or a priori. Then, the dataset is projected onto the interval and an additive noise mechanism is used to ensure privacy. The trimmed mean functions similarly, removing extreme observations instead of projecting them. Existing estimators using these approaches have at least one of the previously mentioned drawbacks. In particular, Tsfadia et al., [34] considers approximately differentially private mean estimation. For stronger forms of privacy, many mean estimation algorithms require certain components of the population distribution as algorithm inputs, such as the number of moments possessed by the distribution [18, 23] or a lower bound on the variance of the population distribution [9]. On the other hand, algorithms aimed at being more “practical”, in the sense that they can be implemented without knowledge of the population distribution, lack minimax optimality or similar statistical guarantees [34, 13]. The most similar work is that of Bun and Steinke, [9], who develop a zero-concentrated differentially private trimmed mean estimate, combined with smooth sensitivity. Their estimator avoids many of the drawbacks listed, however, a lower bound on the variance is required, and it is not clear how to choose the smoothing parameter. In addition, their approach removes data from the estimator, which can be problematic at low sample sizes (see Section 4).
@@
Multivariate estimators can be applied to univariate estimation problems. In the multivariate setting, approximate differentially private mean estimation has been well-studied [25, 5, 17, 12, 6, 11]. Under stronger privacy guarantees, many authors have considered mean estimators for a sample of subgaussian data [22, 7, 10, 4, 26]. Several authors have relaxed this assumption using moment conditions on the population distribution, through contamination models, or both [20, 35]. Unfortunately, some of these algorithms are not currently practically computable [20, 5, 26]. For the estimators that are computable, at small sample sizes ($n<1000$) we have observed empirically that they do not perform as well as univariate estimators applied coordinate-wise, and may be unstable or highly sensitive to tuning parameters (see Section 5). To use these estimators in subsample-and-aggregate, we require thousands of subsamples, and, consequentially, a very large sample size. This may not always be feasible.
@@
Beyond mean estimation, several authors have considered improvements or variants of subsample-and-aggregate [33, 3, 21, 1]. Most of these focus on the “subsample” step of the algorithm [30, 3, 21]. Amin et al., [1] and Smith, [33] consider alternatives for the private aggregator. Smith, [33] show that using a private widened winsorized mean as the aggregator results in the same asymptotic distribution as the nonprivate estimator used in the subsample step. However, this aggregator is sensitive to input bounds, unless we relax the setting to approximate differential privacy. Amin et al., [1] consider an approximately differentially private, multivariate median aggregator in the setting of linear regression.
1.2 Organization and contributions
@@
Our contributions are listed as follows: (i) We develop a new, univariate, differentially private mean estimator, designed to be used as the aggregator in subsample-and-aggregate. (ii) We show that this estimator is minimax optimal for several large classes of distributions, even under adversarial contamination (Theorem 3.1). As a corollary, we present a finite sample deviation bound for subsample-and-aggregate with the new aggregator (Corollary 5.1). (iii) We demonstrate that the estimator performs well relative to competitors in simulation (Section 4). (iv) We demonstrate via real data analysis that common differentially private multivariate mean estimators do not perform well as the aggregator, even with a dataset with 8000 observations (Section 5). (v) As a by-product of our analysis, we introduce a zero-concentrated differentially private version of the unbounded private quantile algorithm [13], as well as some concentration inequalities for these quantile estimates (Lemma D.1 and Lemma E.1E.2). In addition, we correct an error in the proof of Theorem 1 of Lugosi and Mendelson, [28], see Appendix F.
The paper is organized as follows. Section 2 presents the necessary background on differential privacy and mean estimation. Section 3 presents the private modified winsorized mean, our main theoretical result, and a more practical version of the private modified winsorized mean. Section 4 presents a simulation study, comparing the proposed estimator with various existing estimators. Section 5 presents a theoretical and empirical analysis of the improved subsample-and-aggregate algorithm.
@@ 2.1 Problem statement
First, we introduce notation to formalize the problem. For a positive integer $m$, let $[m]=\{1,\ldots,m\}$. Suppose that we have access to a random sample of size $2n$, of which at most $2\eta n$ points, for $0\leq\eta\leq 1$, are corrupted in that they have been changed to arbitrary values. The corrupted sample is denoted $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$. The uncorrupted points, denoted $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, are assumed to be an independent, identically distributed sample from a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$. The aim is to estimate $\mu$ while maintaining differential privacy (to be defined in Section 2.2). This setup is summarized as Condition 1.
@@
The uncorrupted points, $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, are an independent, identically distributed sample from a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$. The points $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ are equal to the uncorrupted points at all but $2\eta n$ points, $0\leq\eta\leq 1$, which have been set to arbitrary, unknown values.
@@ 2.2 Differential privacy
Next, we introduce the essential concepts from differential privacy used in this work [15, 16]. Define a dataset of size $n\in\mathbb{N}$ to be a set of $n$ real numbers, and let $\mathcal{D}_{n}$ be the set of datasets of size $n$. We say that a dataset $\mathbf{x}_{n}\in\mathcal{D}_{n}$, is adjacent to another dataset $\mathbf{y}_{n}\in\mathcal{D}_{n}$ if $\mathbf{x}_{n}$ and $\mathbf{y}_{n}$ differ by exactly one point. Let $\mathcal{A}_{n}$ be the set of pairs of adjacent datasets of size $n$. Next, let $G_{\mathbf{x}_{n}}$ be a probability distribution over $\mathbb{R}^{d}$ with $d\in\mathbb{N}$ that depends on $\mathbf{x}_{n}$. That is, $G_{.}\colon\mathcal{D}_{n}\to\mathcal{M}_{1}(\mathbb{R}^{d}),$ where $\mathcal{M}_{1}(\mathbb{R}^{d})$ is the set of (Borel) probability distributions on $\mathbb{R}^{d}$. Assuming that $G_{\mathbf{x}_{n}}$ is absolutely continuous, let $g_{\mathbf{x}_{n}}$ be the associated density. We can now define (pure) differential privacy (PDP).
Definition 2.1.
The quantity $\theta\sim G_{\mathbf{x}_{n}}$ is $\varepsilon$-differentially private ($\varepsilon$-PDP) if
$\sup_{(\mathbf{y}_{n},\mathbf{z}_{n})\in\mathcal{A}_{n}}\sup_{x\in\mathbb{R}}\frac{g_{\mathbf{y}_{n}}(x)}{g_{\mathbf{z}_{n}}(x)}\leq e^{\varepsilon}.$
Here, $\theta$ is a $d$-dimensional differentially private quantity and $\varepsilon$ is the privacy budget, where values of $\varepsilon$ closer to zero imply higher levels of privacy. In some settings, pure differential privacy can be too strict. A relaxed version of differential privacy is zero-concentrated differential privacy (zCDP) [8]. For $\alpha>1$, let ${\rm D}_{\alpha}(F|G)$ denote the $\alpha$-Renyi Divergence between two distributions, $F$ and $G$.
Definition 2.2.
Given $\rho>0$, the quantity $\theta\sim F_{\mathbf{x}_{n}}$ is $\rho$-zero-concentrated differentially private ($\rho$-zCDP) if, for all $\alpha\in(1,\infty)$,
@@
Similar to $\epsilon$ in PDP, $\rho$ is the privacy budget for zCDP, where values of $\rho$ closer to zero imply higher levels of privacy. Quantities that are $\varepsilon$-PDP are also $\varepsilon^{2}/2$-zCDP quantities. The following two properties are useful for developing differentially private estimates.
Proposition 2.1.
@@
Proposition 2.1 first states that functions of differentially private quantities are still private. Additionally, it states that releasing two differentially private quantities results in a jointly differentially private estimate, with a larger privacy parameter.
Next, we introduce the simplest way to generate a differentially private statistic: additive noise mechanisms. For a univariate statistic $T\colon\mathcal{D}_{n}\to\mathbb{R}$, define the global sensitivity of $T$ as
@@
Proposition 2.2.
Let $T\colon\mathcal{D}_{n}\to\mathbb{R}$ be any statistic with $\mathrm{GS}(T)<\infty$. If $Z$ is a standard Laplace random variable, then
$T(\mathbf{x}_{n})+\mathrm{GS}(T)Z/\varepsilon$
is $\varepsilon$-differentially private. This is referred to as the Laplace mechanism. If $Z$ is a standard Gaussian random variable, then
$T(\mathbf{x}_{n})+\mathrm{GS}(T)Z/\sqrt{2\rho}$
is $\rho$-zCDP. This is referred to as the Gaussian mechanism.
Proposition 2.2 illustrates how correctly calibrated noise can be added to a statistic to ensure differential privacy.
@@ 2.3 Lugosi and Mendelson’s modified winsorized mean
We now review the non-private inspiration for the proposed estimator. Lugosi and Mendelson, [28] recently showed that, out of the class of non-private estimators, a modified winsorized mean estimator is minimax optimal, for the model given in Condition 1. Given a trimming proportion $20<p<1/$ 2, the idea is to first use half of the sample, $\mathbf{Y}_{n}^{\prime}=\{Y_{i}^{\prime}\}_{i=1}^{n}$, to estimate the $p$th quantile and the $(1-p)$th quantile of $F$. We denote these quantile estimates by $\hat{\xi}_{p}$ and $\hat{\xi}_{1-p}$, respectively. Then, we project the other half of the observations $\mathbf{X}_{n}^{\prime}=\{X_{1}^{\prime},\ldots,X_{n}^{\prime}\}$ onto the interval $[\hat{\xi}_{p},\hat{\xi}_{1-p}]$. The sample mean of the projected observations is taken as the final estimate of $\mu$. Formally, for $0<q\leq 1$, let $\hat{\xi}_{q}$ be the (left-continuous) $q$th quantile of the empirical distribution defined by $\mathbf{Y}_{n}^{\prime}$. For $a\leq b$, define
@@
The non-private trimmed mean is
@@
Lugosi and Mendelson, [28] prove an upper bound on $|\hat{\mu}_{p}-\mu|$ that holds with probability $1-\delta$, with $\delta$ at most exponentially small in $n$. Their result shows that taking $p=8\eta+12\log(4/\delta)/n$ yields an estimator that is minimax optimal. Here, we extend $\hat{\mu}_{p}$ to the differentially private setting. The motivation is that, for a carefully chosen $p$, the optimality in the nonprivate setting should translate to the private setting, provided we are able to estimate extreme quantiles well under the restriction of differential privacy.

### v3

@@ Abstract
We develop a univariate, differentially private mean estimator, called the private modified winsorized mean, designed to be used as the aggregator in subsample-and-aggregate. We demonstrate, via real data analysis, that common differentially private multivariate mean estimators may not perform well as the aggregator, even in large datasets, motivating our developments. We show that the modified winsorized mean is minimax optimal for several, large classes of distributions, even under adversarial contamination. We also demonstrate that, empirically, the private modified winsorized mean performs well compared to other private mean estimates. We consider the modified winsorized mean as the aggregator in subsample-and-aggregate, deriving a finite sample deviations bound for a subsample-and-aggregate estimate generated with the new aggregator. This result yields two important insights: (i) the optimal choice of subsamples depends on the bias of the estimator computed on the subsamples, and (ii) the rate of convergence of the subsample-and-aggregate estimator depends on the robustness of the estimator computed on the subsamples. Key words and phrases: Differential privacy, robust statistics, mean estimation, subsample-and-aggregate, adversarial contamination.
@@ 1 Introduction
Subsample-and-aggregate is a powerful algorithm that can produce a differentially privately version of any statistic (Nissim et al.,, 2007). For instance, through subsample-and-aggregate, complex machine learning models can be made private easily, with minimal additional programming. The algorithm is a two-step procedure, relying on a “subsample” step and an “aggregate” step. The idea, in brief, is to first split the data into disjoint subsets. Then, on each of these subsets, a nonprivate version of the desired statistic is computed. Next, the set of nonprivate statistics are combined, using a differentially private aggregator, producing a private estimate of the parameter of interest. Generally, if the mean of the subsamples is desired, the private aggregator is a basic differentially private clipped mean (Nissim et al.,, 2007) or a widened winsorized mean (Smith,, 2011). However, to achieve pure differential privacy, both of these aggregators require tight input bounds on the nonprivate statistics.
It is natural to consider more recent, differentially private mean estimators as the private aggregator. These have a weaker dependence, if any, on input bounds. However, currently, replacing the aggregator with any of the existing differentially private mean estimators comes with at least one of the following drawbacks: (i) only a weaker notion of differential privacy is satisfied, (ii) many subsamples are required for a stable mean estimate, resulting in large sample size requirements, (iii) the new algorithm has strong dependence on tuning parameters, (iv) strong assumptions on, or knowledge of, the population distribution is required, or (v) the estimator is not practically computable.
Here, we develop a univariate, differentially private mean estimator, called the private modified winsorized mean, or, PMW mean. Our estimator is designed to be used as the aggregator in subsample-and-aggregate and avoids these drawbacks. The PMW mean can satisfy either pure or zero-concentrated differential privacy, performs well in small samples or under adversarial contamination, has weak dependence on input bounds and other tuning parameters, and can be used to estimate the mean of a general distribution, with only two moments. The proposed estimator combines the winsorized mean of Lugosi and Mendelson, (2021) and the private quantile estimation procedure of Durfee, (2024).
@@ 1.1 Related work
Differentially private mean estimation has been well-studied. In the univariate setting, clipped and trimmed means have been the main approach (Foygel Barber and Duchi,, 2014; Karwa and Vadhan,, 2017; Bun and Steinke,, 2019; Kamath et al.,, 2020; Tsfadia et al.,, 2022; Durfee,, 2024). The clipped mean procedure operates by selecting an interval $[a,b]$, either empirically or a priori, and projecting the dataset onto the interval. An additive noise mechanism is then used to ensure privacy. The trimmed mean functions similarly, removing extreme observations instead of projecting them. Existing estimators using these approaches have at least one of the previously mentioned drawbacks. In particular, Tsfadia et al., (2022) consider approximately differentially private mean estimation. For stronger forms of privacy, many mean estimation algorithms require certain components of the population distribution as algorithm inputs, such as the number of moments (Foygel Barber and Duchi,, 2014; Kamath et al.,, 2020) or a lower bound on the variance(Bun and Steinke,, 2019). Conversely, algorithms aimed that they can be implemented without knowledge of the population distribution lack minimax optimality or similar statistical guarantees (Tsfadia et al.,, 2022; Durfee,, 2024). The most similar work is that of Bun and Steinke, (2019), who develop a zero-concentrated differentially private trimmed mean estimate, combined with smooth sensitivity. Their estimator avoids many of the drawbacks listed, however, a lower bound on the variance is required, and it is unclear how to choose the smoothing parameter. In addition, their approach removes data from the estimator, which can be problematic at low sample sizes (see Section 5).
@@
Multivariate estimators can be applied to univariate estimation problems. In the multivariate setting, approximate differentially private mean estimation has been well-studied (Liu et al., 2021a, ; Brown et al.,, 2021; Esfandiari et al.,, 2022; Duchi et al.,, 2023; Brown et al.,, 2023; Dagan et al.,, 2024). Under stronger privacy guarantees, many authors have considered mean estimators for a sample of subgaussian data (Kamath et al.,, 2018; Bun et al.,, 2019; Cai et al.,, 2021; Biswas et al.,, 2020; Liu et al., 2021b, ). Several authors have relaxed this assumption using moment conditions on the population distribution, through contamination models, or both (Hopkins et al.,, 2022; Yu et al.,, 2024). Unfortunately, some of these algorithms are not currently practically computable (Hopkins et al.,, 2022; Brown et al.,, 2021; Liu et al., 2021b, ). For the computable estimators, at small sample sizes ($n<1000$) we have observed empirically that they do not perform as well as univariate estimators applied coordinate-wise, and may be unstable or highly sensitive to tuning parameters (see Section 6). To use these estimators in subsample-and-aggregate, we would then require thousands of subsamples, and, thus, a very large sample size.
@@
Beyond mean estimation, several authors have considered variants of subsample-and-aggregate (Smith,, 2011; Bassily et al.,, 2018; Jordon et al.,, 2019; Amin et al.,, 2022). Most of these focus on the “subsample” step of the algorithm (Papernot et al.,, 2016; Bassily et al.,, 2018; Jordon et al.,, 2019). Amin et al., (2022) and Smith, (2011) consider alternatives for the private aggregator. Smith, (2011) show that using a private widened winsorized mean as the aggregator results in the same asymptotic distribution as the nonprivate estimator used in the subsample step. However, this aggregator is sensitive to input bounds, unless we relax the setting to approximate differential privacy. Amin et al., (2022) consider an approximately differentially private, multivariate median aggregator in the setting of linear regression.
1.2 Contributions
@@
Our main contributions are: (i) We develop a new, univariate, differentially private mean estimator, designed to be used as the aggregator in subsample-and-aggregate. (ii) We show that this estimator is minimax optimal for several large classes of distributions, even under adversarial contamination (Theorem 1). As a corollary, we present a finite sample deviation bound for subsample-and-aggregate with the new aggregator (Corollary 1). (iii) We demonstrate that the estimator performs well relative to competitors in simulation (Section 5). (iv) We demonstrate via, data analysis, that common differentially private multivariate mean estimators do not perform well as the aggregator, even with data with 2300 subjects (Section 6). (v) As a by-product of our analysis, we introduce a zero-concentrated differentially private version of the unbounded private quantile algorithm (Durfee,, 2024), as well as concentration inequalities for these quantile estimates (Lemma 1 and Lemma 23). In addition, we correct an error in the proof of Theorem 1 of Lugosi and Mendelson, (2021) (Appendices G and C).
@@ 2.1 Problem statement
First, we formalize the problem. Suppose that we have a random sample of size $2n$, with at most $2\eta n$ points, ($20\leq\eta\leq 1/$ 2), corrupted in that they have been changed to arbitrary values, where the corrupted values may depend on the uncorrupted values. Denote the corrupted sample $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$. Assume that the uncorrupted points, denoted $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, are independent and identically distributed according to a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$. The aim is to estimate $\mu$ while maintaining differential privacy (defined in Section 2.2). This setup is summarized as Condition 1.
@@
Take an independent, identically distributed sample $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, drawn from a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$, to be the uncorrupted points. The points $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ then differ arbitrarily from the uncorrupted points at at most $2\eta n$ points, $20\leq\eta\leq 1/$ 2, where the corrupted values may depend on the uncorrupted values.
@@ 2.2 Differential privacy
Next, we introduce the concepts from differential privacy used in this work (Dwork et al.,, 2006; Dwork and Roth,, 2014). Define a dataset of size $n\in\mathbb{N}$ to be a set of $n$ real numbers, and let $\mathcal{D}_{n}$ be the set of datasets of size $n$. We say that a dataset $\mathbf{x}_{n}\in\mathcal{D}_{n}$, is adjacent to another dataset $\mathbf{y}_{n}\in\mathcal{D}_{n}$ if $\mathbf{x}_{n}$ and $\mathbf{y}_{n}$ differ by exactly one point. Let $\mathcal{A}_{n}$ be the set of pairs of adjacent datasets of size $n$. Let $G_{\mathbf{x}_{n}}$ be a probability distribution over $\mathbb{R}^{d}$ with $d\in\mathbb{N}$ that depends on $\mathbf{x}_{n}$. That is, $G_{.}\colon\mathcal{D}_{n}\to\mathcal{M}_{1}(\mathbb{R}^{d}),$ where $\mathcal{M}_{1}(\mathbb{R}^{d})$ is the set of (Borel) probability distributions on $\mathbb{R}^{d}$. If $G_{\mathbf{x}_{n}}$ is absolutely continuous, let $g_{\mathbf{x}_{n}}$ be the associated density. For $\alpha>1$, denote the $\alpha$-Rényi Divergence between distributions, $F$ and $H$, ${\rm D}_{\alpha}(F|H)$. We now define (pure) differential privacy (PDP) (Dwork et al.,, 2006) and zero-concentrated differential privacy (zCDP) Bun and Steinke, (2016).
Definition 1.
The quantity $\theta\sim G_{\mathbf{x}_{n}}$ is $\varepsilon$-differentially private ($\varepsilon$-PDP), $\varepsilon>0$, if
$\sup_{(\mathbf{y}_{n},\mathbf{z}_{n})\in\mathcal{A}_{n}}\sup_{x\in\mathbb{R}}\frac{g_{\mathbf{y}_{n}}(x)}{g_{\mathbf{z}_{n}}(x)}\leq e^{\varepsilon},$
and is $\rho$-zero-concentrated differentially private ($\rho$-zCDP), $\rho>0$, if, for all $\alpha\in(1,\infty)$,
@@
Here, $\theta$ is a $d$-dimensional differentially private quantity and $\varepsilon,\rho$ are the privacy budgets, with smaller values of $\varepsilon,\rho$ producing higher levels of privacy. Quantities that are $\varepsilon$-PDP are also $\varepsilon^{2}/2$-zCDP. The following two properties are useful for developing differentially private estimates.
Proposition 1.
@@
Proposition 1 first states that functions of differentially private quantities are still private, and second that releasing two differentially private quantities results in a jointly differentially private estimate, with a larger privacy parameter.
Next, we introduce a simple way to generate a differentially private statistic: additive noise mechanisms. For a univariate statistic $T\colon\mathcal{D}_{n}\to\mathbb{R}$, define the global sensitivity of $T$ as
@@
Proposition 2.
Let $T\colon\mathcal{D}_{n}\to\mathbb{R}$ be any statistic with $\mathrm{GS}(T)<\infty$. If $Z$ is a standard Laplace random variable, and $Z^{\prime}$ is a standard Gaussian random variable, then
$T(\mathbf{x}_{n})+\mathrm{GS}(T)Z/\varepsilon,\ \text{ and }\ T(\mathbf{x}_{n})+\mathrm{GS}(T)Z^{\prime}/\sqrt{2\rho},$
satisfy $\varepsilon$-PDP and $\rho$-zCDP, respectively.
Proposition 2 illustrates how correctly calibrated noise can be added to a statistic to ensure differential privacy.
@@ 2.3 Lugosi and Mendelson’s modified winsorized mean
We now review the non-private inspiration for the proposed estimator. Lugosi and Mendelson, (2021) recently showed that, out of the class of non-private estimators, a modified winsorized mean estimator is minimax optimal, for the model given in Condition 1. Given a clipping proportion $20<p<1/$ 2, the idea is to first use half of the sample, $\mathbf{Y}_{n}^{\prime}=\{Y_{i}^{\prime}\}_{i=1}^{n}$, to estimate the $p$th quantile and the $(1-p)$th quantile of $F$. We denote these quantile estimates by $\hat{\xi}_{p}$ and $\hat{\xi}_{1-p}$, respectively. Then, we project the other half of the observations $\mathbf{X}_{n}^{\prime}=\{X_{i}^{\prime}\}_{i=1}^{n}$ onto the interval $[\hat{\xi}_{p},\hat{\xi}_{1-p}]$. The sample mean of the projected observations is taken as the final estimate of $\mu$. Formally, for $0<q\leq 1$, let $\hat{\xi}_{q}$ be the (left-continuous) $q$th quantile of the empirical distribution defined by $\mathbf{Y}_{n}^{\prime}$. For $a\leq b$, define
@@
The non-private clipped mean is
@@
Lugosi and Mendelson, (2021) proved an upper bound on $|\hat{\mu}_{p}-\mu|$ that holds with probability $1-\delta$, with $\delta$ at most exponentially small in $n$. Their result shows that taking $p=8\eta+12\log(4/\delta)/n$ yields an estimator that is minimax optimal. Here, we extend $\hat{\mu}_{p}$ to the differentially private setting. The motivation is that, for a carefully chosen $p$, the optimality in the nonprivate setting should translate to the private setting, provided we are able to estimate extreme quantiles well under the restriction of differential privacy.